Ahmad and Ushna start off simultaneously from two towns to meet one another. If Ahmad travels 2 km/h faster than Ushna; they would meet in 3 hours. If Ushna travels 1 km/h slower and Ahmad's speed is two-thirds of his previous speed, they would meet in 4 hours. How far apart are the two towns?
Answer by mananth(16946)

Ahmad travels 2 km/h faster than Ushna; they would meet in 3 hours. If Ushna travels 1 km/h slower and Ahmad's speed is two-thirds of his previous speed, they would meet in 4 hours. How far apart are the two towns?
...
Ushna --- speed = x km/h
Ahmad x+2 km/h
..
combined speed = x+x+2 = 2x+2 km/h
..
D= speed * time
D= 3*(2x+2)
...
Ushna (x-1)
Ahmad 2/3(x+2)
combined speed = 2/3(x+2)+(x-1)
combined speed = 2x/3 + 4/3 +x-1
combined speed = 5x/3 +1/3
...
D= 4*(5x/3+1/3)
...
4*(5x/3+1/3)=3(2x+2)
20x/3+4/3 = 6x+6
20x/3-6x=6-4/3
(20x-18x)/3=14/3
2x=14
2x=14
x= 7 km/h Ushna's speed.
...
D= 3(2x+2)
D=3(14+2)
D= 48 km. Distance between two towns
...
m.ananth@hotmail.ca